You are considering investing RM65000 in new equipment. You estimate that the net cash flows will be RM11000 during the first year, but will increase by RM2500 per year the next year and each year thereafter. The equipment is estimated to have a 11-year service life and a net salvage value of RM5400 at that time. Assume MARR of 9%. a.Calculate the annual capital cost CR (ownership cost) for the Format: 2925 equipment b.Determine the equivalent annual savings c.Is this a wise investment? Y/N Format 45983 Format : A
Answer #1
PV of Cost
Particulars Amount Discounting factor PV
Investment $ 65,000.00 1 $          65,000.00
Net salvage value $ (5,400.00) 0.54703424 $          (2,953.98)
Present value $          62,046.02
Annual capital cost = 62046.0151/Annuity fator for 7 years of 9%= 62046.0151/5.032952835 $ 12,327.95
PV of Cost savings
Year Cost savings Discounting factor PV of Cost savings
1 $ 11,000.00 0.91743119 $          10,091.74
2 $ 13,500.00 0.84167999 $          11,362.68
3 $ 16,000.00 0.77218348 $          12,354.94
4 $ 18,500.00 0.70842521 $          13,105.87
5 $ 21,000.00 0.64993139 $          13,648.56
6 $ 23,500.00 0.59626733 $          14,012.28
7 $ 26,000.00 0.54703424 $          14,222.89
Total $         88,798.96
Equivalent annual savings = PV of savings/Annuity factor = 88798.95671/5.032952835 = $ 17,643.51
NPV of the Investment = 88798.95671-62046.0151 $ 26,752.94
Since NPV is positive this is a wise Investment = Y